OpenStudy (anonymous):

I have a initial value problem:dy/dx + y/x= -1, y(1)=1. Find the integrating factor and the solution of the problem. Please help me for the solution.

OpenStudy (unklerhaukus):

\[\frac{\text dy}{\text dx} +\frac1x y= -1\]

OpenStudy (unklerhaukus):

the integrating factor of an equation of the form \[y^\prime+p(x)y=q(x)\]is\[\mu(x)=e^{\int p(x)\text dx}\]

OpenStudy (anonymous):